Please solve by completing the square
x^2 + 15x = -56

Answer by mananth(16949)

x^2 + 15x = -56
..
x^2 + 15x +225/4= -56+(225/4)
..
(x+7.5)^2=1/4
..
Take the square root of both sides
(x+7.5)= +/-(1/2)
x= -7.5-(1/2)= -8
OR
x=-7.5+1/2=-7
...
m.ananth@hotmail.ca